OpenStudy (anonymous):

What is the value of x in the proportion x+1/x+3=15/21 i need help can some one help me with the steps

OpenStudy (anonymous):

you could try cross multiplying \[\frac{ a }{ b } = \frac{ c }{ d } \rightarrow ad = bc\] in this case: \[\frac{ x + 1 }{ x + 3 } = \frac{ 15 }{ 21 } \rightarrow (x+1) * 21 = (x + 3) * 15\]

OpenStudy (anonymous):

so it would be 3x=1x ?

OpenStudy (anonymous):

well we have (x + 1) * 21 = (x + 3) * 15 this is the same as 21x + 21 = 15x + 45

OpenStudy (anonymous):

i have no idea what im doing sorry

OpenStudy (anonymous):

you have to solve this algebra problem for x: 21x + 21 = 15x + 45

OpenStudy (anonymous):

can you do a example problem ?

OpenStudy (anonymous):